# BIBO stability of port-Hamiltonian systems

### v1

@@ 2.1. BIBO stability of distributed pHS
As the pHS considered here feature finite-dimensional input and output spaces, Theorem 3.4 from [21] provides a necessary and sufficient condition for BIBO stability. That is the system is BIBO stable if and only if the transfer function $\mathbf{G}$ is the Laplace transform of a measure of bounded total variation on $[0,\infty)$.
@@
While in theory this allows to check any pHS of this form for BIBO stability, in practice this condition is often not usable. Consider as an example the system from the proof of Proposition 2.5. Using Theorem 12.2.1 from [12] we find that its transfer function is $\mathbf{G}(s)=\frac{1}{e^{s}-1}$. Already for this simple function no nice closed form expression exists for its inverse Laplace transform, preventing us from a simple application of the criterion. As it turns out, this is the case generically, rendering this approach to proofing respectively disproofing BIBO stability to be of little use in practice if looking at a particular pHS. However, we will see that the transfer function and the properties of its inverse Laplace transform can — somewhat surprisingly — still be used to derive sufficient conditions for subclasses of these systems.
@@
Finally, note that – similar as for exponential stability [24] – no condition only in terms of the matrices $P_{1}$ and $\widetilde{W}_{B}$ can fully classify BIBO stability for pHS.
Proposition 2.6.
@@
There exists a BIBO stable $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ and a ~$\widetilde{\mathcal{H}}\in L^{\infty}\left([a,b],\mathbb{K}^{n\times n}\right)$ such that $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ is not BIBO stable.
@@
Change now the Hamiltonian to arrive at the system $\textrm{pH}_{2}\left(\mathbbm{1}_{2},0,\left[\begin{smallmatrix}\frac{1}{2}&0\\ 0&1\end{smallmatrix}\right],\widetilde{W}_{B},\widetilde{W}_{C}\right)$ Let then $u=\left[\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right]$ be any smooth input such that $u_{1}(t)=0$ and $u_{2}(t)=(-1)^{t+1}$ for $t\in\mathbb{N}$. Then one can show that $y_{2}(t)=x_{2}(0,t)=\frac{1}{9}(-1)^{t+1}\left(3t+4-4\left(-\frac{1}{2}\right)^{t+1}\right)$ for $t\in\mathbb{N}$. Thus $\lim_{t\rightarrow\infty}y_{2}(t)=\infty$ and the system cannot be BIBO stable. ∎
@@ 3. An approach to BIBO stability via decomposition of the transfer function
A first approach to deriving sufficient conditions for BIBO stability is based on applying the characterisation from [21] not in an individual case-by-case manner — which as we saw may be prohibitively hard — but instead to a general expression of the transfer function.
@@ 3.1. An expression for the transfer function
We begin by deriving a particular decomposition of the transfer function of a pHS as described above by methods inspired by the constructions carried out in [24] and [32] for the study of exponential respectively asymptotic stability.
For any pHS and $s\in\mathbb{C}$ we have the fundamental solution $\Psi_{\zeta}^{s}\in\mathbb{K}^{n\times n}$ of the associated ODE
(3.1) $v^{\prime}(\zeta)=-P_{1}^{-1}\left(P_{0}(\zeta)-s\mathcal{H}^{-1}(\zeta)\right)v(\zeta).$
on $[a,b]$ with $\Psi_{a}^{s}=\mathbbm{1}_{n}$ as a continuous function $\zeta\mapsto\Psi_{\zeta}^{s}$ due to [32, Cor. A.2].
Proposition 3.1.
The transfer function $\mathbf{G}(s)$ of $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ satisfying Assumption 2.1 is given for $s\in\rho(A)$ by
(3.2) $\begin{split}\mathbf{G}(s)=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\left(\widetilde{W}_{B}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\right)^{-1}.\end{split}$
@@
from which the claim follows by solving for $\mathbf{G}(s)$. ∎
Remark 3.2.
@@
(3.3) $\Psi_{\zeta}^{s}=e^{-P_{1}^{-1}\left(P_{0}-s\mathcal{H}^{-1}\right)(\zeta-a)}.$
@@
(3.4) $\Psi_{\zeta}^{s}=e^{-P_{1}^{-1}\int_{a}^{\zeta}\left(P_{0}-s\mathcal{H}^{-1}(\tau)\right)\,\mathrm{d}\tau}.$
@@ 3.2. Decomposing $\widetilde{W}_{B}$
$E_{+}:=\textrm{lin}\left\{v\in\mathbb{K}^{n}\middle|\exists\lambda>0:P_{1}v=\lambda v\right\},\quad E_{-}:=\textrm{lin}\left\{v\in\mathbb{K}^{n}\middle|\exists\lambda<0:P_{1}v=\lambda v\right\}.$
@@
$P_{1}^{\pm}:=\iota_{\pm}^{*}\left(\pm P_{1}\right)\iota_{\pm}:E_{\pm}\rightarrow E_{\pm},\qquad Q_{\pm}:=\iota_{\pm}\left(P^{\pm}_{1}\right)^{\frac{1}{2}}\iota^{*}_{\pm}:\mathbb{K}^{n}\rightarrow\mathbb{K}^{n},$
@@
For the further study of the transfer function as expressed in Equation (3.2), it will be beneficial if the input matrix $\widetilde{W}_{B}$ can be decomposed in a particular way with respect to the matrices $Q_{+}$ and $Q_{-}$. We believe the following result providing this decomposition to be of interest in its own right.
Proposition 3.3.
@@
    (3.5) $\widetilde{W}_{B}=\begin{pmatrix}JQ_{+}-LQ_{-}&JQ_{-}-LQ_{+}\end{pmatrix}.$
    In that case $J=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}$ and $L=\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ Q_{+}\end{bmatrix}P_{1}^{-1}$.
@@
      (3.6) $\widetilde{W}_{B}=K\begin{pmatrix}Q_{+}-MQ_{-}&Q_{-}-MQ_{+}\end{pmatrix}.$
@@
    Furthermore, in this case $K=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}$ and $M=K^{-1}\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ Q_{+}\end{bmatrix}P_{1}^{-1}$.
@@
    $J$ $:=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}=\left(W_{1}Q_{+}+W_{2}Q_{-}\right)P_{1}^{-1}$
    $L$ $:=\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ Q_{+}\end{bmatrix}P_{1}^{-1}=\left(W_{1}Q_{-}+W_{2}Q_{+}\right)P_{1}^{-1}.$
@@
    Conversely, assume that $\widetilde{W}_{B}$ is given in the form (3.5). Then
    $\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}=\left(JQ_{+}-LQ_{-}\right)Q_{+}P_{1}^{-1}+\left(JQ_{-}-LQ_{+}\right)Q_{-}P_{1}^{-1}=J$
@@
    By Part 1 of the Lemma and the definitions of $P_{\pm}$ and $Q_{\pm}$, clearly the decomposition (3.6) exists if and only if
    $J=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}=\widetilde{W}_{B}\begin{bmatrix}P_{+}\\ P_{-}\end{bmatrix}\left(Q_{+}+Q_{-}\right)P_{1}^{-1}$
    is invertible, giving the equivalence of (a) and (d). Further, as both $Q_{+}+Q_{-}$ and $P_{1}^{-1}$ are invertible, this is the case if and only if (b) holds.
@@
    Finally, the expressions for $K$ and $M$ follow then directly from Part (1).
Remark 3.4.
@@
    Note that statement (c) of Proposition 3.3 could equivalently be formulated as $\textrm{pH}_{n}\left(P_{1},P_{0},\mathbbm{1}_{n},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ satisfying Assumption 2.1.
@@
    A direct corollary of Proposition 3.3 is that a decomposition (3.6) exists in particular if $\widetilde{W}_{B}$ satisfies Equation (2.2), as in this case the semigroup generation property is independent of the Hamiltonian. This particular result was first derived by explicit calculation in [24, Lem. 2.3].
@@
    If a decomposition (3.6) holds true, then by direct calculation we find that
@@
    so that in particular Equation (2.2) is equivalent to $\mathbbm{1}_{n}\geq MM^{\ast}$ respectively $\|M\|_{\ell^{2}\rightarrow\ell^{2}}\leq 1$. This is the second half of [24, Lem. 2.3].
@@
    An example of a pHS where no decomposition (3.6) exists is given by $P_{1}=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$, $P_{0}=0$, $\mathcal{H}(\xi)=\left[\begin{smallmatrix}1+\xi&0\\ 0&1\end{smallmatrix}\right]$ for $\xi\in[0,1]$ and $\widetilde{W}_{B}=\left[\begin{smallmatrix}1&0&-1&0\\ 0&1&0&1\end{smallmatrix}\right]$. Then by [11, Ex. 3.2] the operator $A_{\mathbbm{1}}$ does not generate a $C_{0}$-semigroup and thus by Proposition 3.3 no such decomposition exists.
For the remainder of this subsection we now take the assumption that the equivalent statements from Proposition 3.3 hold true.
Assumption 3.5.
@@
(3.7) $\widetilde{W}_{B}=K\begin{pmatrix}Q_{+}-MQ_{-}&Q_{-}-MQ_{+}\end{pmatrix}.$
We now consider the middle of the three parts from the expression for $\mathbf{G}(s)$ in Equation (3.2) closer. Note that with the decomposition (3.7) we have
@@
Lemma 3.6.
The matrix $Q_{-}+Q_{+}\Psi_{b}^{s}$ is invertible for any $s\in\rho(\widetilde{A})$ where $\widetilde{A}$ is the operator associated to $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\left[\begin{smallmatrix}Q_{+}&Q_{-}\end{smallmatrix}\right],\widetilde{W}_{C}\right)$. In particular there exists $\alpha\in\mathbb{R}$ such that $\mathbb{C}_{\alpha}\subset\rho(\widetilde{A})$ and hence $Q_{-}+Q_{+}\Psi_{b}^{s}$ is invertible for all $s\in\mathbb{C}_{\alpha}$.
@@
Note first that $\widetilde{A}$ indeed satisfies Assumption 2.1, as it generates a $C_{0}$-semigroup by [12, Thm. 7.2.4, Thm. 10.3.1] as
@@
By Proposition 3.1 the transfer function of the system $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\left[\begin{smallmatrix}Q_{+}&Q_{-}\end{smallmatrix}\right],\widetilde{W}_{C}\right)$ can thus be represented on $\rho(\widetilde{A})$ as $\widetilde{\mathbf{G}}(s)=\widetilde{W}_{C}\left[\begin{smallmatrix}\widetilde{\Psi}_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]\left(\left[\begin{smallmatrix}Q_{+}&Q_{-}\end{smallmatrix}\right]\left[\begin{smallmatrix}\widetilde{\Psi}_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]\right)^{-1}$ where $\widetilde{\Psi}_{\zeta}^{s}$ is the fundamental solution associated to that system. Hence in particular $\left[\begin{smallmatrix}Q_{+}&Q_{-}\end{smallmatrix}\right]\left[\begin{smallmatrix}\widetilde{\Psi}_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]=Q_{-}+Q_{+}\widetilde{\Psi}_{b}^{s}$ is invertible. But from the definition of the fundamental solution it is clear that $\widetilde{\Psi}_{\zeta}^{s}=\Psi_{\zeta}^{s}$ thus proving the claim. ∎
Remark 3.7.
Under the additional assumptions that $\mathcal{H}\in C^{1}\left([a,b],\mathbb{K}^{n\times n}\right)$ and $P_{0}=-P_{0}^{*}$, the operator $\widetilde{A}$ even generates an exponentially stable contraction semigroup [12, Lem. 9.1.4]. Then there even exists a negative $\alpha<0$ satisfying the conditions in Lemma 3.6. Thus $Q_{-}+Q_{+}\Psi_{b}^{s}$ and $(Q_{-}+Q_{+}\Psi_{b}^{s})^{-1}$ are then in particular defined on the imaginary axis and the whole positive right half plane.
Definition 3.8.
For $s\in\mathbb{C}_{\alpha}$ with $\alpha$ from Lemma 3.6, define the matrices
$\mathbf{V}(s):=Q_{-}+Q_{+}\Psi_{b}^{s}\qquad\textrm{and}\qquad\mathbf{U}(s):=\left(Q_{+}+Q_{-}\Psi_{b}^{s}\right)\mathbf{V}(s)^{-1}.$
@@
Lemma 3.9.
The transfer function of a pHS satisfying Assumptions 2.1 and 3.5 is given on $s\in\rho(A)\cap\mathbb{C}_{\alpha}$ with $\alpha$ from Lemma 3.6 by $\mathbf{G}(s)=\mathbf{Z}(s)\left(1-M\mathbf{U}(s)\right)^{-1}K^{-1},$ where $\mathbf{Z}(s)=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\mathbf{V}(s)^{-1}$. This holds in particular on $\mathbb{C}_{\widetilde{\alpha}}\subset\rho(A)\cap\mathbb{C}_{\alpha}$ for some $\widetilde{\alpha}\geq\alpha$.
@@
As mentioned before, one problem in deciding the question of BIBO stability using the characterisation from [21] is finding the inverse Laplace transform of $\mathbf{G}$. Even in the case of a constant Hamiltonian where the fundamental solution $\Psi_{\zeta}^{s}$ takes the simple form (3.3) of a matrix exponential, we see that due to the inverse $\left(\widetilde{W}_{B}\left[\begin{smallmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]\right)^{-1}$ it will in general be difficult to find that inverse transform. With the representation from Lemma 3.9 however, this problem can be circumnavigated in some cases and lead to a sufficient condition for BIBO stability.
Lemma 3.10.
@@
    Let $s\in\mathbb{C}_{\widetilde{\alpha}}$ with $\widetilde{\alpha}$ from Lemma 3.9. Then if the series $\sum_{k=0}^{\infty}\left(M\mathbf{U}(s)\right)^{k}$ converges we have that
    (3.8) $\mathbf{G}(s)=\mathbf{Z}(s)\left(\sum_{k=0}^{\infty}\left(M\mathbf{U}(s)\right)^{k}\right)K^{-1}.$
@@
    then $\textrm{pH}_{n}\left(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ is BIBO stable and we have
    (3.9) $\mathcal{L}^{-1}\left\{\mathbf{G}\right\}=\mathcal{L}^{-1}\left\{\mathbf{Z}\right\}\ast\left(\sum_{k=0}^{\infty}\mathcal{L}^{-1}\left\{\left(M\mathbf{U}\right)^{k}\right\}\right)K^{-1}.$
@@
    By the assumptions the right hand side of Equation (3.9) is well-defined as a measure of bounded total variation $h\in\mathcal{M}\left([0,\infty),\mathbb{K}^{n\times n}\right)$.
    Term-by-term Laplace transforming of the right hand side of Equation (3.9) yields the right hand side of Equation (3.8) and the continuity of the Laplace transform implies that this sum indeed converges to $\mathcal{L}\left\{h\right\}$ or equivalently that Equation (3.9) holds. BIBO stability then follows from [21, Thm. 3.4]. ∎

### v2

@@ 2.1. BIBO stability of hyperbolic boundary control systems
As the system class considered here features finite-dimensional input and output spaces, Theorem 3.4 from [30] provides a necessary and sufficient condition for BIBO stability. That is the system is BIBO stable if and only if the transfer function $\mathbf{G}$ is the Laplace transform of a measure of bounded total variation on $0,\infty)$.
@@
While in theory this allows to check any hBCS of this form for BIBO stability, in practice this condition is often not usable. Consider as an example the system from the proof of Proposition 5. Using Theorem 12.2.1 from [21] we find that its transfer function is $\mathbf{G}(s)=\frac{1}{e^{s}-1}$. Already for this simple function no nice closed form expression exists for its inverse Laplace transform, preventing us from a simple application of the criterion. As it turns out, this is the case generically, rendering this approach to proofing respectively disproofing BIBO stability to be of little use in practice if looking at a particular system. However, we will see that the transfer function and the properties of its inverse Laplace transform can – somewhat surprisingly – still be used to derive sufficient conditions for subclasses of these systems.
@@
Finally, note that – similar as for exponential stability [34] – no condition only in terms of the matrices $P_{1}$ and $\widetilde{W}_{B}$ can fully classify BIBO stability for these systems.
Proposition 7.
@@
There exists a BIBO stable $\Sigma_{n}(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C})$ and a $\widetilde{\mathcal{H}}\in L^{\infty}\left([a,b],\mathbb{K}^{n\times n}\right)$ such that $\Sigma_{n}(P_{1},P_{0},\widetilde{\mathcal{H}},\widetilde{W}_{B},\widetilde{W}_{C})$ is not BIBO stable.
@@
Change now the Hamiltonian to arrive at the system $\Sigma_{2}\left(\mathbbm{1}_{2},0,\left[\begin{smallmatrix}\frac{1}{2}&0\\ 0&1\end{smallmatrix}\right],\widetilde{W}_{B},\widetilde{W}_{C}\right)$ Let then $u=\left[\begin{smallmatrix}u_{1}\\ u_{2}\end{smallmatrix}\right]$ be any smooth input such that $u_{1}(t)=0$ and $u_{2}(t)=(-1)^{t+1}$ for $t\in\mathbb{N}$. Then one can show that $y_{2}(t)=x_{2}(0,t)=\frac{1}{9}(-1)^{t+1}\left(3t+4-4\left(-\frac{1}{2}\right)^{t+1}\right)$ for $t\in\mathbb{N}$. Thus $\lim_{t\rightarrow\infty}|y_{2}(t)|=\infty$ and the system cannot be BIBO stable. ∎
@@ 3. An approach to BIBO stability via decomposition of the transfer function
A first approach to deriving sufficient conditions for BIBO stability is based on applying the characterisation from [30, Thm. 3.4] not in an individual case-by-case manner — which as we saw may be prohibitively hard — but instead to a general expression of the transfer function.
@@ 3.1. An expression for the transfer function
We begin by deriving a particular decomposition of the transfer function of a hBCS as described above by methods inspired by the constructions carried out in [34] and [42] for the study of exponential respectively asymptotic stability.
For any such hBCS and $s\in\mathbb{C}$ we have the fundamental solution $\Psi_{\zeta}^{s}\in\mathbb{K}^{n\times n}$ of the associated ODE
(4) $v^{\prime}(\zeta)=-P_{1}^{-1}\left(P_{0}(\zeta)-s\mathcal{H}^{-1}(\zeta)\right)v(\zeta).$
on $[a,b]$ with $\Psi_{a}^{s}=\mathbbm{1}_{n}$ as a continuous function $\zeta\mapsto\Psi_{\zeta}^{s}$ due to [42, Cor. A.2].
Proposition 8.
The transfer function $\mathbf{G}(s)$ of $\Sigma_{n}(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C})$ satisfying Assumption 1 is given for $s\in\rho(A)$ by
(5) $\begin{split}\mathbf{G}(s)=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\left(\widetilde{W}_{B}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\right)^{-1}.\end{split}$
@@
from which the claim follows by solving for $\mathbf{G}(s)$. The invertibility of $\widetilde{W}_{B}\left[\begin{smallmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]$ thereby follows from the existence of a unique $\mathbf{G}(s)$ solving Equation (LABEL:eq:transferFunctionODE). ∎
Remark 9.
@@
(6) $\Psi_{\zeta}^{s}=e^{-P_{1}^{-1}\left(P_{0}-s\mathcal{H}^{-1}\right)(\zeta-a)}.$
@@
(7) $\Psi_{\zeta}^{s}=e^{-P_{1}^{-1}\int_{a}^{\zeta}\left(P_{0}-s\mathcal{H}^{-1}(\tau)\right)\,\mathrm{d}\tau}.$
@@ 3.2. Decomposing $\widetilde{W}_{B}$
$E_{+}:=\left\{v\in\mathbb{K}^{n}\middle|\exists\lambda>0:P_{1}v=\lambda v\right\},\quad E_{-}:=\left\{v\in\mathbb{K}^{n}\middle|\exists\lambda<0:P_{1}v=\lambda v\right\}.$
@@
(8) $P_{1}^{\pm}:=\iota_{\pm}^{*}\left(\pm P_{1}\right)\iota_{\pm}:E_{\pm}\rightarrow E_{\pm},\qquad Q_{\pm}:=\iota_{\pm}\left(P^{\pm}_{1}\right)^{\frac{1}{2}}\iota^{*}_{\pm}:\mathbb{K}^{n}\rightarrow\mathbb{K}^{n},$
@@
For the further study of the transfer function as expressed in Equation (5), it will be beneficial if the input matrix $\widetilde{W}_{B}$ can be decomposed in a particular way with respect to the matrices $Q_{+}$ and $Q_{-}$. We believe the following result providing this decomposition to be of interest in its own right.
Proposition 10.
Let $\Sigma_{n}(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C})$be a hBCS satisfying Assumption 1 and let $P_{\pm}$ and $Q_{\pm}$ be defined as above. Then the following holds:
@@
    (9) $\widetilde{W}_{B}=\begin{pmatrix}JQ_{+}-LQ_{-}&JQ_{-}-LQ_{+}\end{pmatrix}$.
    In that case $J=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ -Q_{-}\end{bmatrix}P_{1}^{-1}$ and $L=\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ -Q_{+}\end{bmatrix}P_{1}^{-1}$.
@@
      (10) $\widetilde{W}_{B}=K\begin{pmatrix}Q_{+}-MQ_{-}&Q_{-}-MQ_{+}\end{pmatrix}$.
@@
    Furthermore, in this case $K=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ -Q_{-}\end{bmatrix}P_{1}^{-1}$ and $M=K^{-1}\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ -Q_{+}\end{bmatrix}P_{1}^{-1}$.
@@
    $J$ $:=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ -Q_{-}\end{bmatrix}P_{1}^{-1}=\left(W_{1}Q_{+}-W_{2}Q_{-}\right)P_{1}^{-1}$
    $L$ $:=\widetilde{W}_{B}\begin{bmatrix}Q_{-}\\ -Q_{+}\end{bmatrix}P_{1}^{-1}=\left(W_{1}Q_{-}-W_{2}Q_{+}\right)P_{1}^{-1}.$
@@
    Conversely, assume that $\widetilde{W}_{B}$ is given in the form (9). Then
    $\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ Q_{-}\end{bmatrix}P_{1}^{-1}=\left(JQ_{+}-LQ_{-}\right)Q_{+}P_{1}^{-1}-\left(JQ_{-}-LQ_{+}\right)Q_{-}P_{1}^{-1}=J$
@@
    By Part 1 of the Lemma and the definitions of $P_{\pm}$ and $Q_{\pm}$, clearly the decomposition (10) exists if and only if
    $J=\widetilde{W}_{B}\begin{bmatrix}Q_{+}\\ -Q_{-}\end{bmatrix}P_{1}^{-1}=\widetilde{W}_{B}\begin{bmatrix}P_{+}\\ P_{-}\end{bmatrix}\left(Q_{+}-Q_{-}\right)P_{1}^{-1}$
    is invertible, giving the equivalence of (a) and (d). Further, as both $Q_{+}-Q_{-}$ and $P_{1}^{-1}$ are invertible, this is the case if and only if (b) holds.
@@
    Finally, the expressions for $K$ and $M$ follow then directly from Part (1). ∎
Remark 11.
@@
    Note that statement (c) of Proposition 10 could equivalently be formulated as the system $\Sigma_{n}(P_{1},P_{0},\mathbbm{1}_{n},\widetilde{W}_{B},\widetilde{W}_{C})$ satisfying Assumption 1(5).
@@
    A direct corollary of Proposition 10 is that a decomposition (10) exists in particular if $\widetilde{W}_{B}$ satisfies Equation (2), as in this case the semigroup generation property is independent of the Hamiltonian. This particular result was first derived by explicit calculation in [34, Lem. 2.3].
@@
    If a decomposition (10) holds true, then by direct calculation we find that
@@
    so that in particular Equation (2) is equivalent to $\mathbbm{1}_{n}\geq MM^{\ast}$ respectively $\|M\|_{\ell^{2}\rightarrow\ell^{2}}\leq 1$. This is the second half of [34, Lem. 2.3].
@@
    An example of a hBCS where no decomposition (10) exists is given by $P_{1}=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$, $P_{0}=0$, $\mathcal{H}(\xi)=\left[\begin{smallmatrix}1+\xi&0\\ 0&1\end{smallmatrix}\right]$ for $\xi\in[0,1]$ and $\widetilde{W}_{B}=\left[\begin{smallmatrix}1&0&-1&0\\ 0&1&0&1\end{smallmatrix}\right]$. Then by [18, Ex. 3.2] the operator $A_{\mathbbm{1}}$ does not generate a $C_{0}$-semigroup and thus by Proposition 10 no such decomposition exists.
For the remainder of this subsection we now take the assumption that the equivalent statements from Proposition 10 hold true.
Assumption 12.
@@
(11) $\widetilde{W}_{B}=K\begin{pmatrix}Q_{+}-MQ_{-}&Q_{-}-MQ_{+}\end{pmatrix}$.
We now consider the middle of the three parts from the expression for $\mathbf{G}(s)$ in Equation (5) closer. Note that with the decomposition (11) we have
@@
Lemma 13.
[…]
@@
Note first that $\widetilde{A}$ indeed satisfies Assumption 1(5), as it generates a $C_{0}$-semigroup by [21, Thm. 7.2.4, Thm. 10.3.1] as
@@
[…]
Remark 14.
Under the additional assumptions that $\mathcal{H}\in C^{1}\left([a,b],\mathbb{K}^{n\times n}\right)$ and $P_{0}=-P_{0}^{*}$, the operator $\widetilde{A}$ even generates an exponentially stable contraction semigroup [21, Lem. 9.1.4]. Then there even exists a negative $\alpha<0$ satisfying the conditions in Lemma 13. Thus $Q_{-}+Q_{+}\Psi_{b}^{s}$ and $(Q_{-}+Q_{+}\Psi_{b}^{s})^{-1}$ are then in particular defined on the imaginary axis and the whole positive right half plane.
Definition 15.
For $s\in\mathbb{C}_{\alpha}$ with $\alpha$ from Lemma 13, define the matrices
$\mathbf{V}(s):=Q_{-}+Q_{+}\Psi_{b}^{s}\qquad\textrm{and}\qquad\mathbf{U}(s):=\left(Q_{+}+Q_{-}\Psi_{b}^{s}\right)\mathbf{V}(s)^{-1},$
where $\Psi_{\xi}^{s}$ is the fundamental solution associated to the hBCS and $Q_{\pm}$ are the projection operators defined in Equation (8).
@@
Lemma 16.
The transfer function of a hBCS satisfying Assumptions 1 and 12 is given on $s\in\rho(A)\cap\mathbb{C}_{\alpha}$ with $\alpha$ from Lemma 13 by $\mathbf{G}(s)=\mathbf{Z}(s)\left(1-M\mathbf{U}(s)\right)^{-1}K^{-1},$ where $\mathbf{Z}(s)=\widetilde{W}_{C}\begin{bmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{bmatrix}\mathbf{V}(s)^{-1}$. This holds in particular on $\mathbb{C}_{\widetilde{\alpha}}\subset\rho(A)\cap\mathbb{C}_{\alpha}$ for some $\widetilde{\alpha}\geq\alpha$.
@@
As mentioned before, one problem in deciding the question of BIBO stability using the characterisation from [30] is finding the inverse Laplace transform of $\mathbf{G}$. Even in the case of a constant Hamiltonian where the fundamental solution $\Psi_{\zeta}^{s}$ takes the simple form (6) of a matrix exponential, we see that due to the inverse $\left(\widetilde{W}_{B}\left[\begin{smallmatrix}\Psi_{b}^{s}\\ \mathbbm{1}_{n}\end{smallmatrix}\right]\right)^{-1}$ it will in general be difficult to find that inverse transform. With the representation from Lemma 16 however, this problem can be circumnavigated in some cases and lead to a sufficient condition for BIBO stability.
Lemma 17.
Let the assumptions of Lemma 16 hold.
@@
    Let $s\in\mathbb{C}_{\widetilde{\alpha}}$ with $\widetilde{\alpha}$ from Lemma 16. Then if the series $\sum_{k=0}^{\infty}\left(M\mathbf{U}(s)\right)^{k}$ converges we have that
    (12) $\mathbf{G}(s)=\mathbf{Z}(s)\left(\sum_{k=0}^{\infty}\left(M\mathbf{U}(s)\right)^{k}\right)K^{-1}.$
@@
    then $\Sigma_{n}(P_{1},P_{0},\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C})$ is BIBO stable and we have
    (13) $\mathcal{L}^{-1}\left\{\mathbf{G}\right\}=\mathcal{L}^{-1}\left\{\mathbf{Z}\right\}\ast\left(\sum_{k=0}^{\infty}\mathcal{L}^{-1}\left\{\left(M\mathbf{U}\right)^{k}\right\}\right)K^{-1}.$
@@
    By the assumptions the right hand side of Equation (13) is well-defined as a measure of bounded total variation $h\in\mathcal{M}\left([0,\infty),\mathbb{K}^{n\times n}\right)$.
    Term-by-term Laplace transforming of the right hand side of Equation (13) yields the right hand side of Equation (12) and the continuity of the Laplace transform implies that this sum indeed converges to $\mathcal{L}\left\{h\right\}$ or equivalently that Equation (13) holds. BIBO stability then follows from [30, Thm. 3.4].